OpenStudy (anonymous):

medal for answer. saidhari bought a toaster oven for $60 dollars. this price was 1/4 off the list price. what is the list price?

OpenStudy (anonymous):

Take the difference between the original and after discount price, letting x = original price. x - x/4 = $60 3x/4 = $60 Then, multiply both sides by 4/3 to get.. 3x/4 * 4/3 = $60 * 4/3 x = $80 I hope this helps!

OpenStudy (anonymous):

so its $80 then do you understand it? :)

OpenStudy (anonymous):

@Chriss1234 are you there?????

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now how do i give you a medal?

OpenStudy (anonymous):

@HELPryry

OpenStudy (anonymous):

thank you